1a) Two trains emit whistles of the same frequency, 410Hz. If one train is at rest and the other is traveling at 100 km/hr away from John at rest, what will John detect as the beat frequency?
Answer #1

Velocity, v = 100 x 5/18 = 27.78 m/s

New frequency, f' = f /(1 + vs/v)

f' = 410 /[1 + (27.78/343)] = 379.28 Hz

Beat frequency, Fb = | 410 - 379.28 | = 30.72 Hz

Comment in case any doubt please rate my answer....